Hello, I tried to solve the equation 5(square root)x + 1 = 3 by doing 5(sq root) x = 3 and (5(sq. rt.) x) ^ 5 = 3^5 and I my answer is x=243. However, I'm not sure so can you help me please. Thank you
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

5(square root)x + 1 = 3
If you mean:
5sqrt(x)+1=3
5sqrtx=3-1
5sqrtx=2
sqrtx=2/5
x=(2/5)^2
x=4/25 ans.
----------------------------
OR do you mean:
5sqrt(x+1)=3
sqrt(x+1)=3/5
x+1=(3/5)^2
x+1=9/25
x=9/25-1
x=(9-25)/25
x=-16/25
x=-4/5 ans.

Answer by Alan3354(69443)

Hello, I tried to solve the equation 5(square root)x + 1 = 3 by doing 5(sq root) x = 3 and (5(sq. rt.) x) ^ 5 = 3^5 and I my answer is x=243
----------------
Do you mean 5*sqrt(x+1) = 3 ?
If so,
Square both sides
25*(x+1) = 9
25x+25 = 9
25x = -16
x = -16/25
If that's not what you meant, you'll have to be more specific.